# Some remarks on the exponential separation and dimension preserving approximation for sets and measures

### v1

@@ 3. Main Results
It is noted that the IFS does not satisfy SSC, but there exists a subcollection of $2$-nd level IFS satisfying SSC (see [17, 35]). In view of this, the previous theorem is a weaker version of the Hochman result [18, Theorem 1.4].
To our knowledge, we could not find any proof or statement of the following lemma in the literature; however, some partial mentions for Dirac measures can be seen in [30] and also, since convolution is the smoothing operation, we believe $\dim_{L^{q}}(\mu*\nu)\geq\dim_{L^{q}}(\mu).$ So, it is crucial to include the complete proof with all details for the reader’s convenience.
@@
Let $\mathcal{S}(\mathbb{R}^{m})=\big{\{}\sum_{j=1}^{n}c_{j}\delta_{x_{j}}:x_{j}\in\mathbb{R}^{m},c_{j}>0,\sum_{j=1}^{n}c_{j}=1,n\in\mathbb{N}\big{\}}$ and $\mu$ be a probability measure with finite support. Then, for a fixed $\omega\in\mathcal{S}(\mathbb{R}^{m}),$
@@
Further, we consider a non-zero compactly supported measure $\lambda\in\mathcal{S}_{\beta}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\lambda^{\prime}=\nu*\lambda_{1},$ where $\lambda_{1}(D):=\lambda(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Lemmas 3.17 and 3.18 implies that $\dim_{L^{q}}(\lambda^{\prime})=\dim_{L^{q}}(\lambda_{1})=\dim_{L^{q}}(\lambda)=\beta,$ that is, $\lambda^{\prime}\in\mathcal{S}_{\beta}.$ In view of the properties of convolution mentioned in Section 2, we get
$d_{L}(\nu,\lambda^{\prime})$ $=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\nu-\int_{\mathbb{R}^{m}}fd\lambda^{\prime}\Big{|}\Big{\}}$
$=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\Big{(}\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big{)}-\int_{\mathbb{R}^{m}}fd(\nu*\lambda_{1})\Big{|}\Big{\}}.$
@@
$\int_{\mathbb{R}^{m}}fd(\nu*\lambda_{1})$ $=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d\nu(x)d\lambda_{1}(y)=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d(\sum_{j=1}^{n}c_{j}\delta_{x_{j}})(x)d\lambda_{1}(y)$
$=\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j}+y)d\lambda_{1}(y).$
@@
$d_{L}(\nu,\lambda^{\prime})$ $=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\sum_{j=1}^{n}c_{j}f(x_{j})-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\lambda_{1}(y)\Big{|}\Big{\}}\hskip 11.38092pt$
$=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j})d\lambda_{1}(y)-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\lambda_{1}(y)\Big{|}\Big{\}}$
$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}|f(x_{j})-f(y+x_{j})|d\lambda_{1}(y)\Big{\}}$
$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}\text{Lip}(f)\|y\|_{2}d\lambda_{1}(y)\Big{\}}.$
We claim that the support of $\lambda_{1}$ is a subset of $\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}.$ Let $C\subset\mathbb{R}^{m}\setminus\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}.$ If $x\in C,$ then $\|x\|_{2}>\frac{\epsilon}{4\sqrt{m}}$ and $\frac{8\sqrt{m}diam(K)\|x\|_{2}}{\epsilon}>2diam(K).$ So, $\lambda_{1}(C)=\lambda(\frac{8\sqrt{m}diam(K)C}{\epsilon})=0.$ The previous estimation is reduced to
$d_{L}(\nu,\lambda^{\prime})$ $\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\int_{\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}}\text{Lip}(f)\|y\|_{2}d\lambda_{1}(y)\Big{\}}$
$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\int_{\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}}\text{Lip}(f)\frac{\epsilon}{2}d\lambda_{1}(y)\Big{\}}\leq\frac{\epsilon}{2}.$
@@
$d_{L}(\mu,\lambda^{\prime})\leq d_{L}(\mu,\nu)+d_{L}(\nu,\lambda^{\prime})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon,$
@@
Definition 3.20.
[22] The Fourier transform of $\mu\in\mathcal{P}(\mathbb{R}^{m})$ is defined as
$\hat{\mu}(\xi)=\int e^{-2\pi i\xi.x}d\mu(x),~~~~\xi\in\mathbb{R}^{m}.$
Recall that the Fourier transform of convolution $\mu*\nu$ has the property $\widehat{\mu*\nu}(\xi)=\hat{\mu}(\xi)\hat{\nu}(\xi),$ for details; see [22]. It is said that $\mu$ is a Rajchman measure if $\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0.$ The Riemann-Lebesgue lemma says that $\mu$ is Rajchman whenever it is absolutely continuous with respect to the Lebesgue measure; refer to [27], and the references cited therein. Solomyak [32, 33] studied the Fourier decay for self-similar measures and homogeneous self-affine measures.
@@
If the Fourier transform of a measure $\nu$ is bounded and $\mu$ is Rajchman (with power Fourier decay), then $\mu*\nu$ is also a Rajchman (with power Fourier decay). For any Rajchman measure $\mu$ and real number $\beta>0$, we have $\lim_{|\xi|\to\infty}|\hat{\nu}(\xi)|=0$, where $\nu(D)=\mu(\beta D)$ for each Borel set $D$.
@@
Since $\nu$ is bounded, $\nu(\mathbb{R}^{m})<\infty.$ Consider
$\lim_{|\xi|\to\infty}|\widehat{\mu*\nu}(\xi)|=\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)\hat{\nu}(\xi)|=\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)||\hat{\nu}(\xi)|=K\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0.$
@@
$\lim_{|\xi|\to\infty}|\hat{\nu}(\xi)|=\lim_{|\xi|\to\infty}\Big{|}\int e^{-2\pi i\xi.x}d\nu(x)\Big{|}=\lim_{|\xi|\to\infty}\Big{|}\int e^{-2\pi i\xi.\beta x}d\mu(x)\Big{|}=0,$
@@
The set $\mathcal{S}_{R}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0\}$ is dense in $\mathcal{P}(\mathbb{R}^{m})$. Moreover, for each $\beta\in[0,m]$, the sets $\mathcal{S}_{R}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0\}$ and $\mathcal{S}_{F}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\hat{\mu}(\xi)\text{ has power Fourier decay}\}$ are dense in $\mathcal{P}(\mathbb{R}^{m})$.
@@
Choose a non-zero compactly supported Rajchman measure $\lambda\in\mathcal{S}_{R}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\lambda^{\prime}=\nu*\lambda_{1},$ where $\lambda_{1}(D):=\lambda(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Proposition 3.21 implies that $\lambda^{\prime}\in\mathcal{S}_{R},$ that is, $\lambda^{\prime}$ is Rajchman. Following the lines of Theorem 3.19, one gets $d_{L}(\mu,\lambda^{\prime})<\epsilon,$ showing that $\mathcal{S}_{R}$ is dense in $\mathcal{P}(\mathbb{R}^{m}).$ Further, to show $\mathcal{S}_{R}^{\beta}$ is dense in $\mathcal{P}(\mathbb{R}^{m}),$ choose $\lambda\in\mathcal{S}_{R}^{\beta}.$ This concludes the proof. ∎
@@ 4. Conclusion & future remarks
Theorem 4.1.
Let $0\leq\alpha\leq m$ and $\mu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, there exist Borel probability measures $\nu,\lambda$ such that $\mu=\nu*\lambda$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\lambda)=\alpha.$
Theorem 4.2.
Let $0\leq\alpha\leq m$, and $\mu$ be a finite Borel measure such that $\dim_{L^{q}}(\mu)\leq\alpha.$ Then, there exist finite Borel measures $\nu,\lambda$ such that $\mu=\nu+\lambda$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\lambda)=\alpha.$

### v2

@@ 3. Main Results
It is noted that the IFS does not satisfy SSC, but there exists a subcollection of $2$-nd level IFS satisfying SSC (see [18, 43]). In view of this, the previous theorem is a weaker version of the Hochman result [19, Theorem 1.4].
The upcoming lemmas are elementary observations on the behaviour of the $L^{q}$ dimension, and their proofs are accommodated for young readers. It is remarkable here that Feng et al. [14, Lemma 2.2(ii)] proved one-sided inequality for general Borel probability measures by defining $L^{q}$ dimension in terms of integration.
@@
Let $\mathcal{S}(\mathbb{R}^{m})=\big\{\sum_{j=1}^{n}c_{j}\delta_{x_{j}}:x_{j}\in\mathbb{R}^{m},c_{j}>0,\sum_{j=1}^{n}c_{j}=1,n\in\mathbb{N}\big\}$ and $\mu$ be a Borel probability measure with finite support. Then, for a fixed $\omega\in\mathcal{S}(\mathbb{R}^{m}),$
@@
The upcoming theorems interestingly explore the approximation aspects of a Borel probability measure with a sequence of Borel probability measures preserving
  • $L^{q}$ dimension.
  • Rajchman measures.
  • $L^{q}$ dimension and Rajchman.
  • $L^{q}$ dimension and Fourier transform with power Fourier decay.
Readers obviously think that, why we are interested in highlighting and studying some specific dense subsets of the space of Borel probability measures concerning the $L^{q}$ dimension and Rajchman or power Fourier decay. The Rajchman character of a measure essentially provides information about local regularity and is closely connected to the dimension theory. The Riemann-Lebesgue lemma says that a measure $\eta\in\mathcal{P}(\mathbb{R}^{m})$ is Rajchman whenever it is absolutely continuous with respect to the Lebesgue measure; refer to [33], and the references cited therein. Thus, if $\eta$ is not Rajchman, then it is singular. As an example, the Bernoulli convolutions, except for the particular set of parameters, satisfy the Rajchman property. If for some constant $C,$ we have $|\hat{\eta}(x)|\leq C|x|^{-\frac{t}{2}},$ for all $x\in\mathbb{R}^{m},$ then any set which supports $\eta$ has the Hausdorff dimension atleast $t.$ In fractal geometry, the classification or identification of a class of measures or IFS satisfying specific properties is always of great interest to researchers. Recently, Bárány and Käenmäki [3] showed that self-similar sets on $\mathbb{R}$ which neither satisfy ESC nor have exact overlaps are uncountable. Olsen [27] constructed a $G_{\delta}$ dense set in the space of the Borel probability measure on a compact subset of $\mathbb{R}^{m}$ with upper $L^{q}$ dimension equal to zero with respect to metric $d_{L}$. Recent researches clearly indicate that we want to explore the space $\mathcal{P}(\mathbb{R}^{m})$ and are curious to know whether almost (or typically) all Borel probability measures are Rajchman, or all self-affine measures are Rajchman, and so on. Rapaport [33] commented on the Rajchman property of self-similar measures on $\mathbb{R}^{m}.$ So, going through all the recent works, we thought, can we get some dense subsets of $\mathcal{P}(\mathbb{R}^{m})$ preserving $L^{q}$ dimension or Rajchman property? Our upcoming theorems answer these queries interestingly. Now, the future remarks what we can comment on all these upcoming dense sets to be $G_{\delta}$ dense.
@@
Further, we consider a non-zero compactly supported measure $\eta\in\mathcal{S}_{\beta}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\eta^{\prime}=\nu*\eta_{1},$ where $\eta_{1}(D):=\eta(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Lemmas 3.16 and 3.17 implies that $\dim_{L^{q}}(\eta^{\prime})=\dim_{L^{q}}(\eta_{1})=\dim_{L^{q}}(\eta)=\beta,$ that is, $\eta^{\prime}\in\mathcal{S}_{\beta}.$ In view of the properties of convolution mentioned in Section 2, we get
$d_{L}(\nu,\eta^{\prime})$ $=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\int_{\mathbb{R}^{m}}fd\nu-\int_{\mathbb{R}^{m}}fd\eta^{\prime}\Big|\Big\}$
$=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\int_{\mathbb{R}^{m}}fd\Big(\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big)-\int_{\mathbb{R}^{m}}fd(\nu*\eta_{1})\Big|\Big\}.$
@@
$\int_{\mathbb{R}^{m}}fd(\nu*\eta_{1})$ $=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d\nu(x)d\eta_{1}(y)=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d(\sum_{j=1}^{n}c_{j}\delta_{x_{j}})(x)d\eta_{1}(y)$
$=\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j}+y)d\eta_{1}(y).$
@@
$d_{L}(\nu,\eta^{\prime})$ $=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\sum_{j=1}^{n}c_{j}f(x_{j})-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\eta_{1}(y)\Big|\Big\}\hskip 11.38092pt$
$=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j})d\eta_{1}(y)-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\eta_{1}(y)\Big|\Big\}$
$\leq\sup_{f\in\text{Lip}_{1}}\Big\{\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}|f(x_{j})-f(y+x_{j})|d\eta_{1}(y)\Big\}$
$\leq\sup_{f\in\text{Lip}_{1}}\Big\{\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}\text{Lip}(f)\|y\|_{2}d\eta_{1}(y)\Big\}.$
We claim that the support of $\eta_{1}$ is a subset of $\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}.$ Let $C\subset\mathbb{R}^{m}\setminus\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}.$ If $x\in C,$ then $\|x\|_{2}>\frac{\epsilon}{4\sqrt{m}}$ and $\frac{8\sqrt{m}diam(K)\|x\|_{2}}{\epsilon}>2diam(K).$ So, $\eta_{1}(C)=\eta(\frac{8\sqrt{m}diam(K)C}{\epsilon})=0.$ The previous estimation is reduced to
$d_{L}(\nu,\eta^{\prime})$ $\leq\sup_{f\in\text{Lip}_{1}}\Big\{\int_{\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}}\text{Lip}(f)\|y\|_{2}d\eta_{1}(y)\Big\}$
$\leq\sup_{f\in\text{Lip}_{1}}\Big\{\int_{\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}}\text{Lip}(f)\frac{\epsilon}{2}d\eta_{1}(y)\Big\}\leq\frac{\epsilon}{2}.$
@@
$d_{L}(\mu,\eta^{\prime})\leq d_{L}(\mu,\nu)+d_{L}(\nu,\eta^{\prime})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon,$
@@
  • (i)
    Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ If $\mu$ is Rajchman (with power Fourier decay), then $\mu*\nu$ is also a Rajchman (with power Fourier decay).
  • (ii)
    Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ be any Rajchman measure. Let $\beta>0$ be any real number and $\eta(D)=\mu(\beta D)$ for each Borel set $D$. Then,
    $\lim_{|\zeta|\to\infty}|\hat{\eta}(\zeta)|=0.$
@@
It is well-known (and obvious) that the Fourier transform of a finite measure is bounded. As $\nu\in\mathcal{P}(\mathbb{R}^{m}),$ there exists $K\in\mathbb{R}$ such that $|\hat{\nu}(\zeta)|\leq K,$ for all $\zeta\in\mathbb{R}^{m}.$ Consider
$\lim_{|\zeta|\to\infty}|\widehat{\mu*\nu}(\zeta)|=\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)\hat{\nu}(\zeta)|=\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)||\hat{\nu}(\zeta)|\leq K\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)|=0.$
@@
$\lim_{|\zeta|\to\infty}|\hat{\eta}(\zeta)|=\lim_{|\zeta|\to\infty}\Big|\int e^{-2\pi i\zeta.x}d\eta(x)\Big|=\lim_{|\zeta|\to\infty}\Big|\int e^{-2\pi i\zeta.\beta x}d\mu(x)\Big|=0,$
@@
The set $\mathcal{S}_{R}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)|=0\}$ is dense in $\mathcal{P}(\mathbb{R}^{m})$. Moreover, for each $\beta\in[0,m]$, the sets $\mathcal{S}_{R}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)|=0\}$ and $\mathcal{S}_{F}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\hat{\mu}(\zeta)\text{ has power Fourier decay}\}$ are dense in $\mathcal{P}(\mathbb{R}^{m})$.
@@
Choose a non-zero compactly supported Rajchman measure $\eta\in\mathcal{S}_{R}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\eta^{\prime}=\nu*\eta_{1},$ where $\eta_{1}(D):=\eta(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Proposition 3.19 implies that $\eta^{\prime}\in\mathcal{S}_{R},$ that is, $\eta^{\prime}$ is Rajchman. Following the lines of Theorem 3.18, one gets $d_{L}(\mu,\eta^{\prime})<\epsilon,$ showing that $\mathcal{S}_{R}$ is dense in $\mathcal{P}(\mathbb{R}^{m}).$ Further, to show $\mathcal{S}_{R}^{\beta}$ is dense in $\mathcal{P}(\mathbb{R}^{m}),$ choose $\eta\in\mathcal{S}_{R}^{\beta}.$ This concludes the proof. ∎
@@ 4. Conclusion & future remarks
Issue 1. Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ and $\alpha\in\mathbb{R}$ with $0\leq\alpha\leq\dim_{L^{q}}(\mu).$ Then, there exist Borel probability measures $\nu,\eta$ such that $\mu=\nu*\eta$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\eta)=\alpha.$
Issue 2. Let $0\leq\alpha\leq m$, and $\mu$ be a finite Borel measure such that $\dim_{L^{q}}(\mu)\leq\alpha.$ Then, there exist finite Borel measures $\nu,\eta$ such that $\mu=\nu+\eta$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\eta)=\alpha.$